# Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures

### v1

@@
In conclusion, this work elucidates a new and straightforward method to engineer designer electronic structures in artificial bulk-like systems of transition metal dichalcogenides. By interfacing two flakes of few-layer TMDs, out-of-plane $k_{z}$-dispersions are generated between the limits of the parent compounds, facilitated by charge transfer where necessary. This produces a scenario in which the bulk-like band dispersions of 3D systems and the layer-locked electron behavior of 2D systems coexist within the same structure. These contrasting behaviors interplay, most notably in the production of $k_{z}$-mediated topological phase transitions that stabilize surface layer-localised topological states with properties set by the local crystal parameters, unlike any topologically non-trivial system existing in nature.
@@
When combined with controlled lattice-mismatching, this approach could offer opportunities to engineer a coexistence of strong electronic correlations and non-trivial band topology that, while extremely rare naturally, are pathways towards exotic phases such as topological superconductivity, quantum spin liquids and topological Mott insulators Tokura (2022). Alternatively the non-equivalence of oppositely localized surface states, enforced by inherent $c$-axis structural asymmetry, could be exploited for 2D transistor-like components where the asymmetric surface electron transport can be controlled with band filling or layer-thickness. While demonstrated using flakes with sizes of only tens of microns, new methodologies for creating wafer-scale 2D material flakes Kang et al. (2017) could be exploited for this purpose. Together, the simplicity of this approach and the enormity of the phase space available together provide a new and accessible route towards a function-led construction of three-dimensional artificial systems.
Acknowledgements
@@
O. J. C., M. S. F. and M. T. E. knowledge funding support from ARC Discovery Project DP200101345. M. T. E. acknowledges funding support from ARC Future Fellowship FT2201000290. O. J. C., D. M. and M.T.E. acknowledge travel funding provided by the International Synchrotron Access Program (ISAP) managed by the Australian Synchrotron, part of ANSTO, and funded by the Australian Government. D. M. was supported by the FLEET Centre of Excellence (ARC grant CE170100039). A.A., R.V.B. and M.S.B. gratefully acknowledge the Research Infrastructures at the Center for Computational Materials Science at the Institute for Materials Research for allocations on the MASAMUNE-IMR supercomputer system (Project No. 202112-SCKXX-0510) and MAHAMERU BRIN HPC facility under the National Research and Innovation Agency of Indonesia. M.S.B. acknowledges support from Leverhulme Trust (Grant No. RPG-2023-253). A.A. had support from the Indonesia Endowment Fund for Education (LPDP), NIB/202209223311735. R.V.B. and M.S.B. are grateful to E-IMR center at the Institute for Materials Research, Tohoku University, for continuous support. B. C. and S. L. H. acknowledge the facilities, and the scientific and technical assistance of Microscopy Australia (ROR: 042mm0k03) enabled by NCRIS and the Government of South Australia at Flinders Microscopy and Microanalysis (ROR: 04z91ja70), Flinders University (ROR: 01kpzv902). The work at the University of Warwick was supported by EPSRC, UK through Grant EP/T005963/1. This research used resources of the Advanced Light Source, which is a DOE Office of Science User Facility under contract no. DE-AC02-05CH11231. This work was performed in part at the Melbourne Centre for Nanofabrication (MCN) in the Victorian Node of the Australian National Fabrication Facility (ANFF)
@@
Author Contributions OJC, DM, T-H-YV, MTHB, FM and MTE performed the nano-ARPES measurements. OJC and BC performed the nano-ESCA measurements. T-H-YV performed atomic force microscopy on the WSe2-NbSe2 heterostructure. AA, RVB and MSB performed the theoretical calculations. CJ, AB and ER provided user support at BL 7 MAESTRO of the Advanced Light Source. BC and SLH provided access for the nano-ESCA instrument at Flinders University, South Australia. SHL and ZM supplied semiconductors used in device construction GB supplied the NbSe2 crystals used for device construction. OJC conceived the project and wrote the manuscript with significant contributions from MSB and MTE and further input from all other authors. OJC was responsible for the overall project planning and direction.
Data Availability
Data are available from the authors upon reasonable request. xxxx
I Supplemental Materials for: Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures
Identifying component TMD flakes for heterostructure construction
[…]
In Sup. Fig. 1(a-b), optical images of the component flakes of (WSe2)3/(MoSe2)5 (a) and (MoSe2)3/(WSe2)3 are presented. These flakes, exfoliated onto PDMS using a standard scotch tape method, have thicknesses determined through comparison to photoluminescence spectra and optical images of reference samples.
Heterostructures assembled from these flakes are constructed in such a way that regions of individual components remain exposed laterally (see e.g. Fig. 2(a) of the main text), allowing for photoemission spectra to be acquired directly under identical experimental conditions to those of the heterostructures. This procedure allows for a direct comparison of band structures between component flakes and stacked heterostructures.
In 2H-TMDs, there is a one-to-one correspondence between the number of distinct quantized $k_{z}$-subbands of the $\Gamma$-centered $d_{z^{2}}$ orbital manifolds and the total thickness in monolayers Splendiani et al. (2010). This provides direct in situ verification of flake thicknesses through the enumeration of these doubly-degenerate bands in ARPES spectra. In Sup. Fig. 1(c, d, g) band dispersions close to the $\Gamma$-K direction are displayed for 5ML-MoSe2, 3ML-MoSe2 and 3ML-WSe2, with 5, 3 and 3 quantized states, respectively, indicated explicitly with dashed lines in Sup. Fig. 1(c), (f) and (i), respectively.
The 3ML-MoSe2 and 3ML-WSe2 spectra are compared to density functional theory calculations (see Methods) along the $\Gamma$-K direction. Agreement is very strong across the energy-momentum range shown.
Finally, in Sup. Fig. 1 (f,i), photon energy dependent spectra are acquired for $k_{\parallel}=0$ for both 3ML-MoSe2 (f) and 3ML-WSe2 (i). The dashed line indicates the $k_{z}$ values where the corresponding in-plane spectra are taken in (d) and (g).
The $k_{z}$ dependent datasets shown here and in Fig. 2 of the main text are acquired by continuously varying the photon energy. The relationship between the photon energy, $h\nu$ and $k_{z}$ is determined using a free-electron final state approximation Damascelli (2004):
$k_{z}=\sqrt{\frac{2m_{e}}{\hbar^{2}}(V_{0}+E_{k}(\cos^{2}(\theta))},$
where the inner potential, $V_{0}$, is taken to be 13 eV for all $k_{z}$ spectra in this work, in line with the value found for photon energy dependent studies on numerous bulk 2H-TMD systems (e.g. Riley et al. (2014)).
The varying spectral weight between the individually non-dispersive $k_{z}$ sub-bands (labeled in (f) and (i)) approximates the bulk $k_{z}$ dispersion of the infinite layer system Le Févre et al. (2024). With increasing layer number, these quantized states form a continuous $k_{z}$ dispersion, and the in-plane band dispersions of the highly three-dimensional bands become significantly $k_{z}$-broadened, thereby appearing diffuse, reflecting the small electron mean free path perpendicular to the surface Damascelli (2004).
Lack of rotational disorder in asymmetric heterostructures
Figure S2: (a-b) K-$\Gamma$-K and K-M-K band dispersions as extracted from a three-dimensional ($k_{x}$, $k_{y}$, $E-E_{\text{F}}$) dataset of the eight-layer 3ML-WSe2/5ML-MoSe2 heterostructure. (c) Constant-energy $k_{x}$-$k_{y}$ contours at the energies indicated in (b). 81 eV photons were used for all panels
In Supplemental Figure 2, constant energy $k_{x}-k_{y}$ contours and extracted high symmetry $\Gamma$-K and K-M-K dispersions are shown for the (WSe2)3/(MoSe2)5 heterostructure. Despite the large relative twist angle in both heterostructures between the WSe2 and MoSe2 flakes (see Sup. Fig. 3 for methodology), there is no evidence for rotational disorder. Note in particular the absence of a second set of the spin-orbit split local $d_{xy}$ orbital-derived VBM from the K points. The observed splitting at K is determined by the local crystal structure from the top WSe2 layer, in line with the $d_{xy}$-orbital character and the discussions surrounding Figure 1 of the main text. There are no band features consistent with that of the lower MoSe2 flake. This precludes the scenario that the observed eight-layer spectrum in Sup. Fig. 2(a) and in Figs. 2 and 4 of the main text derive simply from a superposition of spectra from weakly-interacting top and bottom flakes. Indeed, the electron mean free path is on the order of 5.8  Å  for the 81 eV photons used here, comparable to the thickness of only a single 1H unit of a 2H-TMD. Instead all eight visible sub-bands are present in the experimentally accessible top layers. We note that the quality of the photoemission data in Fig. 2 and Sup. Figs. 1-5 reflect the high quality of our heterostructure devices and exfoliation methods (see Methods), though we note the presence of subtle image warping due to instrumentation limitations, explaining both the differing size of the K point valleys and a slight shadowing of the bands at high $k_{x}$ in Sup. Fig. 2(c).
Twist angle determination between top and bottom flakes
[…]
Supplemental Figure 3 shows three band dispersions from three different regions of the sample housing the (MoSe2)3/(WSe2)3 heterostructure discussed in the main text. Band dispersions and constant energy contours are displayed for three regions selected by moving the sample laterally without rotation: (MoSe2)3 (b,e), (WSe2)3 (c,f) and (MoSe2)3/(WSe2)3(d, g). Data are collected using a nano-ESCA instrument at Flinders University, South Australia, with 21.2 eV photons from a Helium plasma discharge lamp.
Despite the much lower photon energy and elevated measurement temperature (300 K), these spectra are entirely consistent with those presented in the main text. Note also the continued lack of rotational disorder (as discussed for the eight layer heterostructure in Sup. Fig. 2). Indicated in Sup. Fig. 3 (e-f) are the in-plane azimuthal rotations applied to the raw data to achieve the high symmetry alignment in the Figure. The twist angle can be determined simply by calculating the difference between the azimuthal angle for the top and bottom TMD flakes. The twist angle for this sample is large at 17.5 degrees, and therefore band folding associated with the Moiré potential should not be expected in this system. The twist angle for the eight layer WSe2-MoSe2 heterostructure is similarly large.
Supplementary datasets for (WSe2)4/(NbSe2)x
Figure S4: (a) K-$\Gamma$-K and K-M-K and $\Gamma$-M-$\Gamma$ band dispersions as extracted from a three-dimensional ($k_{x}$, $k_{y}$, $E-E_{\text{F}}$) dataset of the six-layer 4ML-WSe2/2ML-NbSe2 heterostructure. (c) Constant-energy $k_{x}$-$k_{y}$ contours at the energies indicated in (b). The Brillouin zone is overlaid in purple, and high symmetry points labeled. 91.5 eV photons were used for all panels.
Sup. Fig. 4 shows constant energy contour $k_{x}$-$k_{y}$ mapping and extracted high symmetry dispersions for the (WSe2)4/(NbSe2)2 region of the sample shown in Fig. 2(i) of the main text.
These data demonstrate how the Fermi pocket created by interfacing the WSe2 and NbSe2 flakes is present in both the first and second Brillouin zones. The constant energy contours also exhibit no rotational disorder and no K valleys consistent with isolated NbSe2, again demonstrating that all observed bands are present within the top layers.
[…]
Supplemental Figure 5 compares the three dimensional electronic structure of (WSe2)4/(NbSe2)2 as shown in Figure 2 of the main text, to equivalent datasets from (WSe2)4 and (WSe2)4/(NbSe2)x, $x>>2$. All spectra are referenced to the Fermi level.
The in-plane band dispersion (b, close to $\Gamma$-K) and $k_{z}$ dispersion (c, taken at $k_{\parallel}$=0) for 4ML-WSe2 is consistent with those found for 3ML-WSe2 in Supplemental Fig. 1, with one extra $k_{z}$ subband in the $d_{z^{2}}$ manifold, consistent with the increased layer number. These spectra are compared to the case where the same 4ML-WSe2 flake overlaps with a bulk NbSe2 crystal. The number of $k_{z}$-subbands is unchanged due to the bulk-nature of the bottom component, consistent with the arguments in the main text. Despite this, there are two notable changes relative to free-standing 4ML-WSe2. Firstly, there are additional band features, most notably the conical bands appearing between $k_{z}$-subbands 3 and 4. In Sup. Fig. 5d, an additional spectra taken with 70.5 eV photons is displayed to increase the spectral weight of this conical feature. The second major change is the shift of all bands to shallower binding energies, with the topmost $k_{z}$-subband forming a sole Fermi pocket. While tangential to the findings concerning intermediate-regime asymmetric heterostructures, this result suggests yet another potential approach towards band structure engineering of few-layer TMD flakes. With a careful choice of substrate, the fermiology and band structures can be significantly modified.
Sup. Fig. 5g-i again consider the (WSe2)4/(NbSe2)2 case discussed in the main text. An additional band dispersion using 70.5eV photons (g) is displayed for comparison to the bulk case (d), and shows how a similar conical feature is formed.
We note that, unlike for the WSe2-MoSe2 heterostructures, one can see that there is a superposition of the stacked region with the top 4ML-WSe2 flake here. This has two possible origins: the probing light spot is at the boundary of the overlap (purple) and WSe2 (red) areas, or the less-than-pristine overlap region, as seen in Sup. Fig. 5(a), precludes full and complete charge transfer between the flakes.
Regardless of the origin, the significant charge transfer in the overlap region ensures that the $d_{z^{2}}$-derived bands remain separated from those originating from the top layer alone, thus facilitating an unambiguous enumeration of bands belong to the stacked heterostructure. We also note that the presence of only one extra copy of the WSe2 K point valleys shows that the charge transfer in this region is binary, and thus the observed sub-bands cannot be explained by a series of superimposed band structures with slightly different chemical potentials.
Orbital-dependent $c-$axis localisation
In Supplemental Figures 6 and 7, full layer-resolved DFT calculations for the K-$\Gamma$-K axis is shown for the (MoSe2)3/(WSe2)3 (Sup. Fig. 6) and the (WSe2)3/(MoSe2)5 (Sup. Fig. 7) heterostructures. Consistent with the discussions in the main text and with reference to Fig. 3, bands deriving entirely from out-of-plane orbitals are unchanged through the heterostructure (though intensities do vary), whereas bands deriving entirely from in-plane orbitals are strongly layer-dependent.
The topological surface states (TSSs) discussed in the main text and shown in Fig. 4 are visible here only in the surface layers and not close to the interface between the two compound types. This demonstrates how the TSSs are a product of the quasi-$k_{z}$ dispersion of the full heterostructure, and do not arise from one or both of the components in isolation.
Figure S6: Layer-integrated and layer-projected density functional calculations for the (MoSe2)3/(WSe2)3 heterostructure
Figure S7: Layer-integrated and layer-projected density functional calculations for the (WSe2)3/(MoSe2)5 heterostructure

### v2

@@ IV CONCLUSIONS
In conclusion, this work elucidates a new and straightforward method to engineer designer electronic structures in artificial bulk-like systems of transition metal dichalcogenides. By interfacing two few-layer TMD flakes, emergent bands with characteristics derived from both parent compounds are created, delocalised throughout the heterostructure and facilitated by charge transfer where necessary. This produces a scenario in which the bulk-like band dispersions of 3D systems and the layer-locked electron behavior of 2D systems coexist within the same asymmetric structure. One particularly striking consequence of these contrasting behaviors is the production of $k_{z}$-mediated topological phases that stabilize surface layer-localised topological states with properties set by local crystal parameters, unlike any topologically non-trivial system existing in nature.
@@
When combined with controlled lattice-mismatching, this approach could offer opportunities to engineer a coexistence of strong electronic correlations and non-trivial band topology that, while extremely rare naturally, are pathways towards exotic phases such as topological superconductivity, quantum spin liquids and topological Mott insulators Tokura (2022). Alternatively the non-equivalence of oppositely localized surface states, enforced by inherent $c$-axis structural asymmetry, could be exploited for 2D transistor-like components where the asymmetric surface electron transport can be controlled with band filling or layer-thickness. While demonstrated using flakes with sizes of only tens of microns, new methodologies for creating wafer-scale 2D material flakes Kang et al. (2017) could be exploited for this purpose. We envision that combinations of any of the >> 30 stable TMD systems Chhowalla et al. (2013) could produce hybrid three-dimensional band properties analogous to those shown here with further applicability to other layered materials, including two-dimensional magnets. Together, the simplicity of this approach and the enormity of the phase space available together provide a new and accessible route towards a function-led construction of mixed-dimensional artificial systems.
V ACKNOWLEDGMENTS
@@
O. J. C. thanks Phil D. C. King for useful discussions. O. J. C., M. S. F. and M. T. E. acknowledge funding support from ARC Discovery Project DP200101345. M. T. E. acknowledges funding support from ARC Future Fellowship FT2201000290. O. J. C., D. M. and M.T.E. acknowledge travel funding provided by the International Synchrotron Access Program (ISAP) managed by the Australian Synchrotron, part of ANSTO, and funded by the Australian Government. D. M. was supported by the FLEET Centre of Excellence (ARC grant CE170100039). A.A., R.V.B. and M.S.B. gratefully acknowledge the Research Infrastructures at the Center for Computational Materials Science at the Institute for Materials Research for allocations on the MASAMUNE-IMR supercomputer system (Project No. 202112-SCKXX-0510) and MAHAMERU BRIN HPC facility under the National Research and Innovation Agency of Indonesia. M.S.B. acknowledges support from Leverhulme Trust (Grant No. RPG-2023-253). A.A. had support from the Indonesia Endowment Fund for Education (LPDP), NIB/202209223311735. R.V.B. and M.S.B. are grateful to E-IMR center at the Institute for Materials Research, Tohoku University, for continuous support. B. C. and S. L. H. acknowledge the facilities, and the scientific and technical assistance of Microscopy Australia (ROR: 042mm0k03) enabled by NCRIS and the Government of South Australia at Flinders Microscopy and Microanalysis (ROR: 04z91ja70), Flinders University (ROR: 01kpzv902). The work at the University of Warwick was supported by EPSRC, UK through Grant EP/T005963/1. This research used resources of the Advanced Light Source, which is a DOE Office of Science User Facility under contract no. DE-AC02-05CH11231. This work was performed in part at the Melbourne Centre for Nanofabrication (MCN) in the Victorian Node of the Australian National Fabrication Facility (ANFF)
@@
Author Contributions OJC, DM, T-H-YV, MTHB, FM and MTE performed the nano-ARPES measurements. OJC and BC performed the nano-ESCA measurements. OJC constructed the heterostructures shown throughout the manuscript, with the exception of the bilayer system shown in Sup. Fig. 4 which was constructed by T-H-YV. T-H-YV also performed atomic force microscopy on the WSe2-NbSe2 heterostructure. AA, RVB and MSB performed the theoretical calculations. CJ, AB and ER provided user support at BL 7 MAESTRO of the Advanced Light Source. BC and SLH provided access for the nano-ESCA instrument at Flinders University, South Australia. SHL and ZM supplied semiconductors used in device construction GB supplied the NbSe2 crystals used for device construction. OJC conceived the project and wrote the manuscript with significant contributions from MSB and MTE and further input from all other authors. OJC was responsible for the overall project planning and direction.